OpenStudy (anonymous):

Solve (sin2A)(tanA) = 1 for 0º≤A≤180º.

OpenStudy (anonymous):

\[\sin2A=2\sin A \cos A \times \frac{sinA}{cosA} = 2\sin^2A]=1\] \[\sin^2A=\frac{1}{2}\] A=?